OpenStudy (anonymous):

For the geometric sequence -8, 16, -32, find a5

OpenStudy (anonymous):

For a geometric sequence the nth term a(n) = a(1)*r^(n-1). In this case we have a(1) = 5 and r = -2, so a(n) = 5*(-2)^(n-1) and thus a(5) = 5*(-2)^(5-1) = 5*16 = 80.

OpenStudy (anonymous):

find r then \[a _{n}=a r ^{n-1}\] where a is first term, r is common ratio.

OpenStudy (anonymous):

does anyone know the answer to this

OpenStudy (texaschic101):

first we need to find the common ratio by dividing the 2nd term by the first term. 16/-8 = -2...and you will notice that every term, when multiplied by -2, gives you the next term. now we will use this formula : an = a1 * r^(n - 1) n = the term you want to find a1 = first term = -8 now we sub a5 = -8 * -2^(5 - 1) a5 = -8 * -2^4 a5 = -8 * -16 a5 = 128

OpenStudy (anonymous):

\[a _{5}=-8*\left( -2 \right)^{5-1}=-8\left( -2 \right)^4=-8*16=-128\]

OpenStudy (texaschic101):

oops...my bad....surjithayer is correct...because -2^4 is 16...not -16....sorry about that